The investment demand and the market for money for a country are shown in the graphs below. Suppose the economy currently has a recessionary gap of $100 billion and the MPC is 0.8. Indicate the money supply that the central bank should target if it wants to bring real GDP to the full-employment level.

 

Instructions: Use the tool provided 'Sm2' to draw the new money supply curve. Use the graph on the right to plot your line such that the first point touches the horizontal axis.
Answer #1

Ans )

In the Given figure on the left it can be illustrated that the current investment demand = $20 at the rate of interest equal to 0.8.

Marginal propensity to consume is equal to 0.8.

We know that the Fiscal multiplier formula is as follows:

Multiplier = 1/(1-MPC)

Multiplier = 1/(1-0.8)

Multiplier = 5

Fiscal multiplier is the automatic stabilizer in an economy in which the government conducts an expansionary spending policy which further leads to the multifold increase in the aggregate output in the economy.

It is given in the question that the recessionary gap is equal to $100.

Aggregate output = ( multiplier ) X ( investment)

We are already given the values of aggregate output and multiplier so we can calculate the value of the investment demand necessary to close the recessionary care as follows:

100 = 5 × investment demand

Investment = 100/5 = 20

The government needs to increase the investment by $20 billion which means that final investment demand must equal $40 million and for that the rate of interest must equal 4%. For this to happen, the money supply must be increased to $200 billion.

New graph of the money supply and demand is depicted below in which the money supply has been increased from $100 billion to $200 billion.
Answer #2

Investment Demand The Market for Money 12% 12% New Money Supply 10% 10% 8% 8% Real rate of interest (percent) 6% Real rate ofmAns. Investment increase required to bring the economy back to full employment level = Recessionary gap / Spending multiplier

Here, Spending multiplier = 1/(1-MPC) = 1/(1-0.8) = 5

Thus, investment increase required = 100 billion /5 = $20 billion

At present the interest rate in the money market is 8% (where money demand and money supply curve interests). At 8% interest, the investment spending is $20 billion and required increase in investment spending is $20 billion. So, total investment must be $40 billion which occurs at 4% interest.

So, to decrease interest rate to 4%, the central bank must increase by $100 billion to $200 billion.

Thus, to fill the recessionary gap, the central bank must increase money supply by $100 billion.

* Please don’t forget to hit the thumbs up button, if you find the answer helpful.

Thank You